OpenStudy (anonymous):

Find the measures of the angles of a triangle if they are in a ratio of 2:2:5. A. 40°, 40°, 100° B. 20°, 20°, 140° C. 60°, 60°, 60° D. 30°, 30°, 120°

OpenStudy (anonymous):

Its option A. 40°, 40°, 100°

OpenStudy (anonymous):

how'd you get thar

OpenStudy (anonymous):

Dividing each angle by 20, simce 20 is the common multiple of given numbers.

OpenStudy (anonymous):

40(20) 40(20) right ? @dpasingh

OpenStudy (anonymous):

no \[\frac{40}{20} : \frac{40}{20} : \frac{100}{20} = 2:2:5\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You are welcome.